Open Market Operations

Assume the required reserve ratio is 40% and the Open Market Committee of the FED sells $500 billion in bonds to the public. Assuming banks give out as many loans as possible, what is the total change in the money supply? What is the total change in Transaction Accounts? If the M1 was originally $9,500 billion, what is the new M1?
Answer #1

(i) An open market sale decreases money supply.

Decrease in money supply ($ billion) = Open market sale / Reserve ratio = 500 / 0.4 = 1250

(ii) An open market sale decreases transaction accounts.

Decrease in transaction accounts ($ billion) = Total decrease in money supply - Open market sale = 1250 - 500 = 750

(iii) New M1 ($ billion) = 9500 - 1250 = 8250